### v1

# Metric-aware LLM inference

Abstract
Large language models (LLMs) have demonstrated strong results on a range of NLP tasks. Typically, outputs are obtained via autoregressive sampling from the LLM’s underlying distribution. We show that this inference strategy can be suboptimal for a range of tasks and associated evaluation metrics. As a remedy, we propose metric aware LLM inference: a decision theoretic approach optimizing for custom metrics at inference time. We report improvements over baselines on academic benchmarks and publicly available models.
1 Introduction
Large language models (LLMs) are currently the most capable models across many NLP tasks (OpenAI et al., 2023; Google and et al., 2023; Touvron et al., 2023; Gemini Team and et al., 2023). Owing to their remarkable few- and zero-shot abilities (Wei et al., 2022; Kojima et al., 2023), many tasks can be addressed without conducting any additional model training on in-domain datasets: instead, users can obtain useful results by querying a pretrained LLM with a suitably crafted input prompt.
Different NLP tasks employ different evaluation metrics to assess the quality of an LLM. One popular metric is the exact match (EM), which penalises any response not exactly equal to that provided in the dataset annotation. This is an analogue of the conventional classification accuracy.
While the EM is an intuitive metric, there are many cases where it is not suitable. For example, metrics such as squared error, mean absolute error or correlation scores are often preferred on tasks where outputs are numerical, or where categories are ordered. Such tasks occur in prominent NLP challenges such as relevance scoring (Cer et al., 2017) and sentiment analysis (Fathony et al., 2017). Further, BLEU (Papineni et al., 2002) and ROUGE (Ganesan, 2018) scores are popular for natural language tasks such as summarization and machine translation, while $F_{1}$ score is used to evaluate reading comprehension (Joshi et al., 2017) and text segmentation (Lukasik et al., 2020). Moreover, as noted in Schaeffer et al. (2023), the EM score may not properly reflect improvements in model performance due to its discontinuity.
Despite the wide variety of evaluation metrics, LLM inference is typically performed in the same manner for every task: namely, one performs auto-regressive sampling from the LLM’s underlying distribution (see §2). While intuitive, such inference notably does not explicitly consider the downstream evaluation metric of interest. This raises a natural question: is there value in adapting the inference procedure to the evaluation metric at hand?
In this paper, we focus on the zero-shot abilities of LLMs on tasks evaluated on different metrics. We make a key observation that choosing the most likely target for the output corresponds to inherently optimizing for the exact match evaluation metric; consequently, this inference strategy is not optimal when the EM is not the metric of interest. As a remedy to this, we propose metric aware LLM inference (MALI), a decision-theoretic approach where a metric of choice is optimized at inference time under the distribution from the LLM. We demonstrate across datasets and models how our approach can bring large gains over choosing the most likely target.
Figure 1: An illustration of the metric aware LLM inference. First, an input example is passed to an LLM (forward pass). We next obtain samples from the distribution over targets conditioned on the input $x$ (sampling from $p(y\mid x)$). Finally, the obtained samples are used for finding the target optimizing a metric ${m}$ (metric aware LLM inference); see Table 1 for specific solutions corresponding to example metrics.
2 When (Naïve) LLM Inference Fails
We begin by describing the problem setting. Fix a finite vocabulary $V$ of tokens (e.g., words in English). Let $D$ denote a distribution over inputs $x\in X\subseteq{V}^{*}$ comprising of strings of tokens and targets sequences $y\in Y$. Let $p(y\mid x)$ denote the conditional distribution over targets given an input.
A special case of this setting is when $Y\subset\mathbb{R}$ corresponds to numeric targets. Here, we assume that each $y\in Y$ has a unique string representation ${\tt str}(y)\in V^{*}$; for example, the integer $1$ has the string encoding "1". In a slight abuse of notation, we use $p(y\mid x)\stackrel{{\scriptstyle\cdot}}{{=}}p({\tt str}(y)\mid x)$ to denote the conditional probability of output $y$ given input $x$.
A language model (LM) takes a string $x$ as input and predicts an output $\hat{y}\in Y$. Typically, the LM first produces a distribution $\hat{p}(\cdot\mid x)$ over targets, from which a prediction is derived via a suitable inference (or decoding) procedure. Perhaps the most common inference strategy is to choose the mode of $\hat{p}(y\mid x)$:
$\hat{y}(x):=\operatorname*{argmax}_{y\in Y}\hat{p}(y\mid x).$ (1)
In practice, one may approximate the mode by employing greedy decoding or beam search, or sampling multiple candidates with a temperature and picking the among them the one with the highest likelihood score (Naseh et al., 2023).
The quality of an LM’s prediction is measured by some evaluation metric $m(y,\hat{y})$, where we assume that higher values are better. While the exact match (EM), given by $m(y,\hat{y})=\mathds{1}({y=\hat{y}})$, is a commonly used evaluation metric, there are a range of other metrics popularly used to evaluate LMs. These include the (negative) squared error $m(y,\hat{y})=-(y-\hat{y})^{2}$ or absolute error $m(y,\hat{y})=-|y-\hat{y}|$ in the case of numerical targets, or text-based metrics such as $F_{1}$ or BLEU.
A natural goal is to then choose the inference strategy $\hat{y}(x)$ to maximize the metric $m$ of interest, i.e., to maximize the following expected utility:
$\mathbb{E}_{(x,y)\sim D}\left[m(y,\hat{y}(x))\right].$ (2)
As we shall see below, for many choices of metric $m(y,\hat{y}(x))$, picking the mode of the predicted distribution (1) can be sub-optimal for (2).
Example 1.
Consider the task of predicting the star rating (on the scale 11115555) associated with a review text. Suppose $m(y,\hat{y})$ is the negative absolute error between true and predicted star ratings. Given the review text ‘‘This keybord is suitable for fast typers’’, suppose the responses and the associated probabilities from an LM are {‘‘1’’: 0.3,  ‘‘2’’: 0.0,  ‘‘3’’: 0.3,  ‘‘4’’: 0.0,  ‘‘5’’: 0.4}. The mode of the predicted probabilities is ‘‘5’’. In contrast, the maximizer of (2) is the median rating ‘‘3’’. We provide real examples for Amazon reviews of learn probability distributions and the label in Figure 2 (Appendix).
Example 2.
Consider a reading comprehension task, where we use the $F_{1}$ score as the evaluation metric $m(y,\hat{y})$, defined by the harmonic mean of $\text{recall}(y,\hat{y})=\frac{|y\cap\hat{y}|}{|y|}$ and $\text{precision}(y,\hat{y})=\frac{|y\cap\hat{y}|}{|\hat{y}|}$. Suppose for the question ‘‘What is the hottest month in the year’’, the responses and associated probability from an LM are {‘‘July’’: 0.25,  ‘‘July 2023’’: 0.23,  ‘‘Month of July’’: 0.24,  ‘‘May’’: 0.28}. The mode of this distribution is ‘‘May’’; whereas the maximizer of (2) is ‘‘July’’.
3 Metric-aware LLM inference
We seek to design decoding strategies that maximize the expected utility in (2).
3.1 Bayes-optimal inference
In an idealized setting, where we have access to the true conditional probabilities $p(\cdot\mid x)$, it is straight-forward to show that the maximizer of (2) is:
$\hat{y}^{*}(x)\in\operatorname*{argmax}_{y^{\prime}\in Y}\,\mathbb{E}_{y\sim p(\cdot\mid x)}\left[m(y,y^{\prime})\right].$ (3)
When $m$ is the EM metric, the optimal inference strategy is $\hat{y}^{*}(x)\in\operatorname*{argmax}_{y\in Y}\,p(y\mid x),$ which is what common approaches such as greedy decoding seek to approximate.
In general, however, the optimal decoding strategy can have a very different form. For example, for evaluation metrics over numerical targets such as the squared error or the absolute error, the optimal inference strategy is to simply take the mean or median of $p(\cdot|x)$ (Bishop, 2006). Table 1 summarizes these optimal rules for different metrics.
3.2 Approximating the optimal solution
In practice, we mimic the Bayes-optimal solution in (3) with two approximations. First, we replace the true conditional distribution ${p}(\cdot\mid x)$ with the LM’s predicted distribution $\hat{p}(\cdot\mid x)$. This is a reasonable approximation when the LM is pre-trained with next-token prediction task and softmax cross-entropy loss: the latter is a strictly proper loss, whose minimizer under an unrestricted hypothesis class is the true conditional distribution $p(y\mid x)$ (Gneiting and Raftery, 2007).
Second, we estimate the expectation in (3) by sampling $K$ outputs from $\hat{p}(\cdot\mid x)$ with a sampling temperature $T$, and then computing:
$\hat{y}(x)\in\operatorname*{argmax}_{y^{\prime}\in Y}\,\sum_{i=1}^{K}m(y_{i},y^{\prime}).$ (4)
As noted in Table 1, the above maximization can be computed in closed-form for some common metrics. In applications where we neither have a closed-form solution, nor can enumerate $S$ feasibly, we may solve (4) approximately by maximizing over a fixed subset $C\subset Y$ of candidate outputs:
$\hat{y}(x)\in\operatorname*{argmax}_{y^{\prime}\in C}\,\sum_{i=1}^{K}m(y_{i},y^{\prime}).$ (5)
The choice of $C$ would depend on the task and metric at hand. For example, one simple choice is for $C$ to be the $K$ sampled outputs, i.e., $C=\{y_{1},\ldots,y_{K}\}$. In Appendix A.1, we described how we choose $C$ for text generation tasks.
The quality of our inference strategy depends on the sampling temperature $T$, the number of samples $K$, and the choice of $C$. Note that sampling can be done efficiently by caching the Transformer activations for the input prefix when generating different targets; the Transformer computation needs to be repeated only for the different sampled targets.
3.3 Post-hoc temperature scaling
As mentioned, when sampling from $\hat{p}(\cdot\mid x)$, it often helps to apply a temperature scaling to the LM logits to control the diversity of the sampled outputs. This is particularly important in our procedure where we wish to approximate expectations over $\hat{p}(\cdot|x)$ using a few samples.
In practice, one may sample from $\hat{p}(\cdot\mid x)$ with temperature $T=1$, and apply temperature scaling in a post-hoc manner by employing a weighted version of the objective in (4):
$\hat{y}(x)\in\operatorname*{argmax}_{y^{\prime}\in Y}\,\sum_{i=1}^{K}\left(\hat{p}(y_{i}\mid x)\right)^{\alpha}\cdot m(y_{i},y^{\prime}),$ (6)
where $\alpha$ can be seen as the temperature scaling parameter. The above summation is a (scaled) estimate of $\mathbb{E}_{y\sim\hat{p}(\cdot\mid x)}\left[\hat{p}(y\mid x)^{\alpha}\cdot m(y,y^{\prime})\right]$. For probabilities $\hat{p}(y_{i}\mid x)\propto\exp(f(x,y_{i}))$ defined by logits $f(x,y_{i})$, this is equivalent to computing the expectation under the temperature-scaled distribution $\hat{p}_{\alpha}(y\mid x)\propto\exp((1+\alpha)\cdot f(x,y))$, albeit a normalization factor. We consider an analogous weighting scheme for the plug-in estimators of the closed-form solutions reported in Table 1.
Problem Label space $Y$ Metric Optimal decision rule
Classification $1,\ldots,K$ $\mathds{1}({y=\hat{y}})$ $\hat{y}(x):=\operatorname*{argmax}_{y}[p(y\mid x)]$
Regression $\mathbb{R}$ $-(y-\hat{y})^{2}$ $\hat{y}(x):=\mathbb{E}_{y\sim p(\cdot\mid x)}[y]$
Ordinal regression $1,\ldots,K$ $-|y-\hat{y}|$ $\hat{y}(x):=\text{median}[p(\cdot\mid x)]$
Question answering $V^{*}$ $F_{1}(y,\hat{y})$ $\hat{y}(x):=\operatorname*{argmax}_{y^{\prime}}\,\mathbb{E}_{y\sim p(\cdot\mid x)}\left[m(y,y^{\prime})\right]$
Bi-partite ranking ±$\pm 1$ AUC $\hat{y}(x):=p(y=+1|x)$
Multi-partite ranking $1,\ldots,K$ $\text{AUC}^{\text{multi}}$  with costs $c_{y,y^{\prime}}=|y-y^{\prime}|$ $\hat{y}(x):=\mathbb{E}_{y\sim p(\cdot\mid x)}[y]$
Table 1: Optimal inference for different evaluation metrics. For the first three, we have a closed-form solution. For $F_{1}$, we approximate the optimal solution by optimizing over a subset of candidates $C\subset Y$3.2). See (§3.5) for definitions of AUC. For the problems in rows 1–4, the LLM predicts an output in label space $Y$; for the ranking problems in rows 5–6, the LLM predicts a real-valued score, which is then used to rank the samples.
model size greedy decode MALI
argmax mean $*{}^{*}$
STSB XXS 1.078 1.448 1.007 1.028
S 0.685 1.019 0.641 0.649
L 0.628 0.989 0.580 0.610
argmax median $*{}^{*}$
Amazon reviews XXS 0.495 0.826 0.493 0.474
S 0.301 0.444 0.299 0.285
L 0.294 0.541 0.298 0.291
argmax $F_{1}$ ${F_{1}}^{*}$
Trivia-QA XXS 0.314 0.178 0.307 0.302
S 0.620 0.636 0.650 0.678
L 0.886 0.887 0.880 0.886
Table 2: Root mean squared error (RMSE) on STSB dataset (lower is better), Mean absolute error (MAE) on Amazon reviews dataset (lower is better), and $F_{1}$ metrics on Trivia-QA dataset (higher is better) from PaLM-2 models of varying size. We sample $16$ targets for the argmax and MALI methods. All methods rely on $T=1$ temperature for sampling, and the weighted approaches correspond to an effective temperature $T=\frac{1}{4}$. See Tables 5, 6, 7 in the Appendix for results over more temperatures and evaluation metrics (cost weighted multi-class AUC and Pearson correlation).
3.4 Efficiency
Note that sampling can be done efficiently by caching the Transformer activations for the input prefix when generating different targets. The Transformer computation needs to be repeatedly done only for the different sampled targets.
In practice, when the prefix is long compared to the generated targets (e.g. for scoring tasks, where the target score can be just a few tokens length, whereas the prefix can be long as it contains the input text), a forward pass for the prefix tends to take most of the compute time.
Moreover, we generate multiple samples simultaneously, and so, we do not incur a higher cost from generating multiple targets.
3.5 Ranking metrics
Apart from the settings where evaluation metrics are decomposable, i.e., evaluation metrics are a result of aggregation of per-example results from evaluation, there are also metrics which are non-decomposable. An example for such a metric are ranking metrics, where the goal is to correctly order examples, as opposed to predict correct scores or classes. Ranking metrics are appropriate for many important tasks, including ads click-through rate prediction or recommender systems. Below we review different types of ranking metrics and the corresponding optimal decision rules.
Bi-partite ranking
In bi-partite ranking, the goal is to order examples from two classes (and so the label set $Y={\pm 1}$) in such a way that examples from the positive class are above the examples from the negative class. In this case The LLM predicts real valued scores, and so the target space need not be the same as the label space. One common metric used for evaluating bi-partite ranking is the area under the ROC curve (AUC), which is proportional to the number of wrongly ordered pairs of examples in the ranking (Cortes and Mohri, 2003):
$\text{AUC}(\hat{y})\,=\,1-\mathbb{E}\Big{[}\mathds{1}(\hat{y}(x)<\hat{y}(x^{\prime}))\,+\,\frac{1}{2}\cdot\mathds{1}(\hat{y}(x)=\hat{y}(x^{\prime})\,\Big{|}\,y=+1,y^{\prime}=-1\Big{]}$
It can be shown that the optimal decision rule for AUC takes the form of (any strictly monotone transformation of) $\hat{y}^{*}(x)=p(y=+1|x)$ (Clémençon et al., 2008; Uematsu and Lee, 2015).
Multi-partite ranking
Multi-partite ranking is an extension to the bi-partite ranking, where the ranking is over examples from multiple categories. Several multi-category extensions of binary AUC have been considered, for example by counting wrongly ordered pairs of examples in the ranking according to the category indices and weighing them according to predetermined costs (Uematsu and Lee, 2015):
$\text{AUC}^{\text{multi}}(\hat{y})\,=\,1-\mathbb{E}\Big{[}c_{y,y^{\prime}}\cdot\Big{(}\mathds{1}(\hat{y}(x)<\hat{y}(x^{\prime}))\,+\,\frac{1}{2}\cdot\mathds{1}(\hat{y}(x)=\hat{y}(x^{\prime})\Big{)}\,\Big{|}\,y>y^{\prime}\Big{]}$
Uematsu and Lee show that when the costs are equal to the difference between category indices, i.e., $c_{y,y^{\prime}}=|y-y^{\prime}|$, an optimal solution is to use mean distribution over scores: $\hat{y}^{*}(x)=\mathbb{E}_{y\sim p(\cdot|x)}\left[y\right]$. We report results on the cost weighted multi-class AUC in Table 6 in Appendix.
4 Experiments and discussion
We compare our metric-aware inference approach against standard inference strategies on NLP tasks with different evaluation metrics.
Datasets. We summarize the datasets below, and list the prompts used in Table 4 (Appendix).
  1. (i)
    STSB: The Semantic Textual Similarity Benchmark (Cer et al., 2017) comprises of sentence pairs human-annotated with a similarity score from 0 to 5. As it is a regression task, we consider evaluation with root mean squared error.
  2. (ii)
    US Amazon reviews: We evaluate on a sample of $1500$ examples from the US Amazon reviews dataset. From the review text we aim to predict the 5-star rating corresponding to a product review (Ni et al., 2019). Since the task is in the form of ordinal regression, we use mean absolute error as the metric (Fathony et al., 2017).
  3. (iii)
    TriviaQA is a reading comprehension dataset comprising (question, answer, reference text) triplets, where a question is answered against a refence text (Joshi et al., 2017). The evaluation is the $F_{1}$ score (see Example 2).
We use a sample of $1500$ examples for evaluation.
Models. We consider two instruction-tuned model families: FLAN-T5 (Chung et al., 2022) and PaLM-2 (Google and et al., 2023). We report results across different model sizes and temperatures. Unless otherwise stated, we fix the number of samples to $K=16$, and the top-$k$ parameter in decoding to $40$ (Fan et al., 2018)
Methods. We evaluate the following methods (i) greedy decoding, (ii) sampling $K$ candidates and picking the one with the maximum likelihood (argmax), (iii) the proposed MALI approach on the same $K$ samples, and (iv) the temperature scaled variant of MALI in §3.3 (denoted by a ‘*’). For (iv), we choose $\alpha$ so that the effective temperature is $\frac{1}{4}$. For the tasks involving the squared and absolute error metrics, we use the closed-form decision rules in Table 1. For the $F_{1}$ metric, we solve (5) over a candidate set $C$, which we choose to contain the $K$ samples and additional targets derived from them (details in Appendix A.2).
Metric-aware inference helps. In Table 2, we report results across datasets and model sizes. We notice that MALI improves over baselines across all model sizes on STSB and Amazon reviews (for model size S, we see that median performs very similarly to the most likely generated sample). On Trivia QA, we find that on XXS and S, optimizing for $F_{1}$ helps, and for L, simply using the most likely generated sample works most effectively.
We see diminishing improvements with larger models for Trivia QA and Amazon datasets. This coincides with a lowering entropy in predictions with increasing model size for Trivia QA and Amazon datasets, whereas STSB does not exhibit this phenomenon as strongly: see Table 8 (Appendix) for empirical entropy estimates.
We also evaluate MALI with varying number of samples $K$. As seen in Table 9 (Appendix), on STSB with temperature $\frac{1}{4}$, even with as few as two samples, MALI starts to show improvements over greedy decoding.
Sampling versus scoring. So far, when estimating the prediction maximizing the (2), we have used sampling from the LM distribution (see §3.2). Alternatively, if the targets are from a narrow interval (e.g., on the STSB dataset, the values are in the interval $[0,5]$), one can find the probabilities for targets at fixed intervals (e.g. $0,0.5,1.0,\ldots,5.0$), and thus find the solutions from Table 1 and more generally from (3).
In Table 3, we report results from FLAN-T5 on the STSB dataset for both scoring and sampling based estimates for MALI, where the scoring based estimator is based on $11$ equally spaced targets. We find that both sampling and scoring lead to MALI improving over choosing the most likely prediction. Further, we note that sampling is a more effective strategy than scoring the equally spaced targets.
model greedy scoring sampling
FLAN-T5 S 4.419 2.407 2.275
FLAN-T5 L 0.455 0.410 0.373
FLAN-T5 XL 0.508 0.549 0.457
Table 3: Squared error (SE) on STSB from FLAN-T5 models of varying size. Temperature $0.5$ was used for generating targets for the sampling approach.
Conclusions In this work we show how utilizing the output distribution modeled by LLMs in the form of our MALI methods can bring improvements across NLP tasks. Our work also points at importance of further understanding and improving of the quality of model output distribution and calibration, which we found crucial for the MALI methods to work well.
5 Limitations
There are multiple limitations of our work. First, we evaluate our proposed methods on multiple text datasets with numerical and text targets, however, many more types of outputs can be considered, including time series targets, long text targets etc. Next, it would be interesting to more systematically analyze how to solve the objective from ((6)) over many samples for text outputs for metrics like $F_{1}$ or BLEU, e.g. by means of dynamic programming. We also note that the datasets considered in this work are restricted to English. It would be interesting to expand the explorations to datasets in other languages.
6 Ethics Statement
All datasets used in this work are publicly available. No additional user data was collected or released as part of this work. All models used are publicly available and already pretrained, and no finetuning was conducted for any experiments. Instead, all experiments relied on running inference experiments with the models over several thousands of examples. Thus, the CO-2 footprint of this paper is minimal. We do not foresee any significant risks associated with this paper other than improving performance on tasks which are harmful.
7 Acknowledgements
We are thankful to Ziwei Ji for the helpful feedback.
Appendix A Appendix
A.1 Candidate set construction
One simple choice for the candidate set $C$ could be take the $K$ sampled outputs, i.e., $C=\{y_{1},\ldots,y_{K}\}$. One may additionally include in this set transformations on each $y_{i}$ $y_{i}$ or new candidates formed from combining two or more of the samples.
For example, in applications such as reading comprehension or question-answering, where the output is list of keywords that constitute an answer to a question, one often uses a metric such as F1 that balances between precision and recall. In such cases, after sampling, one may additionally include samples formed by concatenating pairs of sampled outputs, i.e., ${\tt concat}(y_{i},{\tt delim},y_{j}),\forall i\neq j$. These concatenated answers have the effect of increasing recall, at the cost of lower precision. We follow that procedure for the Trivia-QA experiments in §4.
A.2 Additional details
In Table 4 we report the prompts we used in our experiments for zero-shot inference.
For all datasets, we use validation splits, and where not available, we use the first $1500$ examples from the train split.
Dataset Prompt
STSB
What is the sentence similarity between the following two sentences measured on a scale of 0 to 5: {Sentence #1}, {Sentence #2}. The similarity measured on a scale of 0 to 5 with 0 being unrelated and 5 being related is equal to
Amazon reviews
What is the rating corresponding to the following review in the scale of 1 to 5, where 1 means negative, and 5 means positive? Only give a number from 1 to 5 with no text. Review: {Review} Rating:
Trivia-QA
Answer the following question without any additional text. Question: {Question}. Answer:
Table 4: Prompts used for different datasets. Curly braces denote inputs specific to an input example.
A.3 Additional experiments
In Table 8 we report empirical entropy estimates as measured based on the $16$ samples generated from the model. We find that entropy decreases as model size increases. We observe a particularly sharp decrease in entropy for the Amazon reviews and Trivia-QA datasets, where for larger model sizes we don’t find improvements from MALI approaches.
In Table 5 we report RMSE on STSB dataset, MAE on Amazon reviews dataset, and $F_{1}$ metrics on Trivia-QA dataset from PaLM-2 models of varying size across multiple temperature values. We find improvements over baselines on STSB and Amazon reviews datasets for most temperatures. For Trivia-QA, we find improvements for XXS and S models for some temperatures, and for L, we don’t find a difference from our methods due to low entropy in the responses (see Table 8). In Table 7 we additionally report Pearson correlation metrics on STSB, confirming the results of MALI improving over autoregressive inference. Lastly, in Table 6 we report cost weighted multi-class AUC with costs corresponding to the difference between the annotated labels: $|y_{1}-y_{2}|$. We find on both STSB and Amazon reviews datasets that the optimal decision rule (mean over the distribution) improves over the baselines.
In Table 9, we report the impact of the number of samples on the results. We note that there is an improvement in the results with the increase in the number of samples, however beyond 8 samples there is a diminishing improvement in practice.
In Figure 2 we report examples from the Amazon dataset and the corresponding: human annotations and samples from the model. Notice how samples cover significant proportions of the ratings. We find that the samples end up in the vicinity of the human annotation, and thus in many cases taking a mean over samples helps improve the prediction over the mode.
(a) It is a nice color of black and my husband likes how it feels in his hand.
(b) This item is a good idea. However, Unless the ear canal is reasonably deep (…) it’s of no use. The plastic hooks that come with it are hard and too small (…). Might be good for children.
(c) One of the sides is made for apple products, the other is just standard usb. Both will work with apple products, just one side (the A side) charges faster. Other than that, it’s fantastic. :D
Figure 2: Examples from the Amazon dataset and the corresponding: human annotations and samples from the model. We find that in many cases, taking mean over model samples is closer to the annotation than taking the mode of the distribution.
model size greedy decode T=$0.25$ T=$0.5$ T=$1.0$
argmax mean w-mean argmax mean w-mean argmax mean w-mean
STSB XXS 1.078 1.126 1.043 1.028 1.241 1.021 0.992 1.448 1.007 0.978
S 0.685 0.787 0.643 0.649 0.908 0.636 0.642 1.019 0.641 0.641
L 0.628 0.729 0.592 0.610 0.852 0.582 0.586 0.989 0.580 0.580
T=$0.25$ T=$0.5$ T=$1.0$
argmax median w-median argmax median w-median argmax median w-median
Amazon reviews XXS 0.495 0.509 0.484 0.474 0.624 0.485 0.487 0.826 0.493 0.493
S 0.301 0.290 0.297 0.285 0.329 0.300 0.297 0.444 0.299 0.299
L 0.294 0.318 0.293 0.291 0.380 0.294 0.293 0.541 0.298 0.295
T=$0.25$ T=$0.5$ T=$1.0$
argmax $F_{1}$ w-$F_{1}$ argmax $F_{1}$ w-$F_{1}$ argmax $F_{1}$ w-$F_{1}$
Trivia-QA XXS 0.314 0.300 0.319 0.318 0.255 0.323 0.326 0.178 0.307 0.304
S 0.620 0.656 0.626 0.678 0.658 0.641 0.662 0.636 0.650 0.650
L 0.886 0.888 0.886 0.888 0.888 0.883 0.887 0.887 0.880 0.885
Table 5: Root mean squared error (RMSE) on STSB dataset (the lower the better), Mean absolute error (MAE) on Amazon reviews dataset (the lower the better), and $F_{1}$ metrics on Trivia-QA dataset (the higher the better) from PaLM-2 models of varying size. We report different methods of inference across different temperatures. For the weighted approaches, we fix the sampling temperature to $T=1$ and accordingly vary the $\alpha$ in (6) so as to arrive at the effective temperature equal to the value reported.
model size greedy decode T=$0.25$ T=$0.5$ T=$1.0$
argmax mean argmax mean argmax mean
STSB XXS $0.797$ $0.755$ $0.797$ $0.755$ $0.797$ $0.755$ $0.797$
XS $0.895$ $0.870$ $0.895$ $0.870$ $0.895$ $0.870$ $0.895$
L $0.905$ $0.885$ $0.905$ $0.885$ $0.905$ $0.885$ $0.905$
Amazon reviews XXS $0.87$ $0.894$ $0.925$ $0.866$ $0.94$ $0.788$ $0.942$
XS $0.9$ $0.91$ $0.925$ $0.914$ $0.941$ $0.9$ $0.958$
L $0.925$ $0.922$ $0.951$ $0.906$ $0.962$ $0.837$ $0.964$
Table 6: Cost-weighted multi-partite AUC metrics on STSB and Amazon datasets (the higher the better). MALI methods improve over the baselines. See §3.5 for the definition of AUC we use. We assume costs to correspond to the difference between the annotated labels: $|y_{1}-y_{2}|$.
model greedy decode T=$0.25$ T=$0.5$ T=$1.0$
argmax mean argmax mean argmax mean
PaLM-2 XXS $0.767$ 0.738 $\mathbf{0.790}$ 0.670 $\mathbf{0.790}$ 0.544 $0.786$
PaLM-2 XS $0.898$ 0.878 $\mathbf{0.915}$ 0.852 $0.913$ 0.821 $0.910$
PaLM-2 L $0.909$ 0.893 $0.920$ 0.881 $0.922$ 0.860 $\mathbf{0.923}$
Table 7: Pearson correlation metrics on STSB. MALI methods improve over the baselines.
model STSB Amazon Trivia-QA
PaLM-2 XXS $1.141$ $1.064$ $1.328$
PaLM-2 XS $1.055$ $0.753$ $0.475$
PaLM-2 L $0.976$ $0.361$ $0.186$
Table 8: Empirical entropy across model sizes and datasets.
samples XXS S L
(Greedy Decode) 1.078 0.685 0.628
2 1.044 0.679 0.624
4 1.036 0.669 0.613
6 1.031 0.664 0.607
8 1.028 0.660 0.603
10 1.025 0.657 0.601
12 1.024 0.655 0.600
14 1.022 0.653 0.599
16 1.021 0.652 0.598
Table 9: RMSE as a function of the number of samples on STSB across PaLM-2 models of varying size. Results for temperature $T=0.25$.
A.4 Related work
LLMs for numerical tasks.
Generative models have been successfuly applied to number prediction, where a number is generated token by token in an autoregressive manner. For example, Gruver et al. [2023] considered it in a zero-shot learning setup for time series prediction, and Liu and Low [2023], Yang et al. [2023] considered the autoregressive finetuning over numerical targets applied to arithmetic tasks. The importance of tokenizing the numerical targets into individual digits has been raised by previous works [Liu and Low, 2023, Yang et al., 2023].
Finetuning approaches for target tasks alignment.
Previous works considered approaches for aligning the models for target datasets. For example, soft prompts were finetuning on target datasets without loosing generalization to other tasks [Wang et al., 2023], and general finetuning was conducted on carefully tailored datasets for improved model robustness [Li et al., 2023]. In our work, we focus on zero-shot setting where no fine-tuning is conducted.
Finetuning approaches for numerical tasks.
Autoregressive finetuning of LLMs on numerical tasks with CoT has been found effective [Liu and Low, 2023]. One line of work for modeling predictive tasks with pre-trained Transformer based models is to add a regression head on top of the transformed/pooled encoded input tokens and finetune the resulting model on numerical targets using a regression loss. This is an approach which has been for encoder based models (e.g. Bert), and has also been applied to encoder-decoder (e.g. T5) models [Liu et al., 2022], and these approaches could be extended to decoder models too. In this work, we focus on the zero shot approaches, and so we leave training approaches for future work.

### v3

@@
-Metric-aware LLM inference
+Regression-aware Inference with LLMs
 Abstract
-Large language models (LLMs) have demonstrated strong results on a range of NLP tasks.
-Typically, outputs are obtained via autoregressive sampling from the LLM’s underlying distribution.
-We show that this inference strategy can be suboptimal for a range of tasks and associated evaluation metrics.
-As a remedy, we propose metric aware LLM inference: a decision theoretic approach optimizing for custom metrics at inference time.
-We report improvements over baselines on academic benchmarks and publicly available models.
+Large language models (LLMs) have shown strong results on a range of applications, including regression and scoring tasks.
+Typically, one obtains outputs from an LLM via autoregressive sampling from the model’s output distribution.
+We show that this inference strategy can be sub-optimal for common regression and scoring evaluation metrics.
+As a remedy, we build on prior work on Minimum Bayes Risk decoding, and propose alternate inference strategies that estimate the Bayes-optimal solution for regression and scoring metrics in closed-form from sampled responses.
+We show that our proposal significantly improves over baselines across datasets and models.
 1 Introduction
-Large language models (LLMs) are currently the most capable models across many NLP tasks (OpenAI et al., 2023; Google and et al., 2023; Touvron et al., 2023; Gemini Team and et al., 2023).
-Owing to their remarkable few- and zero-shot abilities (Wei et al., 2022; Kojima et al., 2023), many tasks can be addressed without conducting any additional model training on in-domain datasets: instead, users can obtain useful results by querying a pretrained LLM with a suitably crafted input prompt.
-Different NLP tasks employ different evaluation metrics to assess the quality of an LLM.
-One popular metric is the exact match (EM), which penalises any response not exactly equal to that provided in the dataset annotation.
+Large language models (LLMs) are currently the most capable models across many NLP tasks (OpenAI et al., 2023; Anil et al., 2023; Touvron et al., 2023; Gemini Team et al., 2024).
+Owing to their remarkable few- and zero-shot abilities (Wei et al., 2022; Kojima et al., 2023), pre-trained LLMs are often applied without any additional training on domain-specific datasets: instead, one may query the LLM with a suitably crafted input prompt.
+More recently, LLMs have been successfuly applied to regression and scoring tasks.
+For example, Gruver et al.
+(2023) explored zero-shot learning for time series prediction; Vacareanu et al.
+(2024) showed how LLMs are remarkably strong at in-context learning for regression tasks; Liu and Low (2023); Yang et al.
+(2023) considered autoregressive fine-tuning over numerical targets applied to arithmetic tasks; and Qin et al.
+(2023) applied LLMs for listwise ranking.
+The quality of an LLM is often assessed using an application-specific evaluation metric.
+One popular metric is the exact match (EM), which penalises any response not exactly equal to the one in the dataset annotation.
 This is an analogue of the conventional classification accuracy.
-While the EM is an intuitive metric, there are many cases where it is not suitable.
-For example, metrics such as squared error, mean absolute error or correlation scores are often preferred on tasks where outputs are numerical, or where categories are ordered.
-Such tasks occur in prominent NLP challenges such as relevance scoring (Cer et al., 2017) and sentiment analysis (Fathony et al., 2017).
-Further, BLEU (Papineni et al., 2002) and ROUGE (Ganesan, 2018) scores are popular for natural language tasks such as summarization and machine translation, while $F_{1}$ score is used to evaluate reading comprehension (Joshi et al., 2017) and text segmentation (Lukasik et al., 2020).
-Moreover, as noted in Schaeffer et al.
-(2023), the EM score may not properly reflect improvements in model performance due to its discontinuity.
+While EM is an intuitive metric, there are many applications where it is not suitable.
+This includes tasks such relevance scoring (Cer et al., 2017) and sentiment analysis (Fathony et al., 2017), where the outputs are numerical or ordinal categories.
+In these cases, one instead prefers metrics such as the squared error, absolute error or ranking scores that take the outputs’ ordinal nature into account.
 Despite the wide variety of evaluation metrics, LLM inference is typically performed in the same manner for every task: namely, one performs auto-regressive sampling from the LLM’s underlying distribution (see §2).
-While intuitive, such inference notably does not explicitly consider the downstream evaluation metric of interest.
-This raises a natural question: is there value in adapting the inference procedure to the evaluation metric at hand?
-In this paper, we focus on the zero-shot abilities of LLMs on tasks evaluated on different metrics.
-We make a key observation that choosing the most likely target for the output corresponds to inherently optimizing for the exact match evaluation metric; consequently, this inference strategy is not optimal when the EM is not the metric of interest.
-As a remedy to this, we propose metric aware LLM inference (MALI), a decision-theoretic approach where a metric of choice is optimized at inference time under the distribution from the LLM.
-We demonstrate across datasets and models how our approach can bring large gains over choosing the most likely target.
-Figure 1: An illustration of the metric aware LLM inference.
-First, an input example is passed to an LLM (forward pass).
-We next obtain samples from the distribution over targets conditioned on the input $x$ (sampling from $p(y\mid x)$).
-Finally, the obtained samples are used for finding the target optimizing a metric ${m}$ (metric aware LLM inference); see Table 1 for specific solutions corresponding to example metrics.
-2 When (Naïve) LLM Inference Fails
-We begin by describing the problem setting.
-Fix a finite vocabulary $V$ of tokens (e.g., words in English).
-Let $D$ denote a distribution over inputs $x\in X\subseteq{V}^{*}$ comprising of strings of tokens and targets sequences $y\in Y$.
-Let $p(y\mid x)$ denote the conditional distribution over targets given an input.
-A special case of this setting is when $Y\subset\mathbb{R}$ corresponds to numeric targets.
+While intuitive, such inference does not explicitly consider the downstream evaluation metric of interest.
+This raises a natural question: is there value in adapting the inference procedure to the evaluation metric at hand for regression and scoring tasks?
+A prominent line of work takes a decision-theoretic approach to the above problem.
+Dubbed as Minimum Bayes Risk (MBR) decoding, this approach seeks to optimize at inference time the metric of choice under the model’s distribution (Bickel and Doksum, 1977; Kumar and Byrne, 2004; Eikema and Aziz, 2020; Bertsch et al., 2023).
+Much of the work on MBR is focused on evaluation metrics for machine translation and text generation tasks, such as the BLEU score (Papineni et al., 2002).
+Of particular interest in this literature are self-consistency based decoding strategies that take a (weighted) majority vote of sampled responses (Wang et al., 2023a), which have shown to provide quality gains in arithmetic and reasoning problems.
+In this paper, we build on the existing literature on MBR to design metric-aware inference strategies for general regression and scoring tasks.
+We first observe that choosing the most likely target for an input corresponds to inherently optimizing for the EM metric, and is consequently not optimal when EM is not the metric of choice.
+As a remedy, we propose estimating the Bayes-optimal output for a metric under the model’s distribution (see Figure 1 for an illustration of our method); we show that this admits a closed-form solution for common regression and ranking metrics, and only requires estimating a simple statistic from the sampled responses.
+In contrast, prior MBR methods for translation and summarization often require heuristically solving an intractable maximization problem (Ehling et al., 2007; Bertsch et al., 2023).
+We show across datasets and models how our approach yields gains over choosing the most likely target, and over self-consistency based approaches.
+Figure 1: Illustration of metric-aware LLM inference for regression and scoring tasks.
+An input $x$ is passed to the LLM, and samples are drawn from the distribution over targets $y$ conditioned on $x$.
+These are then used to find the target optimizing a metric ${m}$ through a closed-form decision rule $\Phi$ (e.g., mean or median); Table 1 presents specific solutions across metrics.
+Algorithm 1 RAIL: Regression-aware Inference with LLMs
+1:input: Model $M$, #samples $K$, sampling temperature $T$, effective temperature $T^{\prime}$, decision rule $\Phi$
+2:for $i=1$ to $K$ do
+3:     $\hat{y}_{i},\hat{p}_{i}=M.\text{\tt generate}(T)$$\triangleright$ $M$.generate returns the sampled target and its probability.
+4:end for
+5:$\alpha=\frac{T}{T^{\prime}}-1$$\triangleright$ Compute post-hoc temperature scaling so that the effective temperature used is $T^{\prime}$
+6:return $\Phi(\hat{y}_{1},\ldots,\hat{y}_{K},\hat{p}_{1},\ldots,\hat{p}_{K},\alpha)$
+2 When (naïve) LLM inference fails on regression tasks
+We begin with the problem setting.
+For a finite vocabulary $V$ of tokens (e.g., words in English), let $D$ denote a distribution over inputs $x\in X\subseteq{V}^{*}$ comprising of strings of tokens, and targets $y\in Y$.
+Let $p(y\,|\,x)$ denote the conditional distribution over targets given an input.
+We consider a special case of this setting where $Y\subset\mathbb{R}$ corresponds to numeric targets.
 Here, we assume that each $y\in Y$ has a unique string representation ${\tt str}(y)\in V^{*}$; for example, the integer $1$ has the string encoding "1".
-In a slight abuse of notation, we use $p(y\mid x)\stackrel{{\scriptstyle\cdot}}{{=}}p({\tt str}(y)\mid x)$ to denote the conditional probability of output $y$ given input $x$.
-A language model (LM) takes a string $x$ as input and predicts an output $\hat{y}\in Y$.
-Typically, the LM first produces a distribution $\hat{p}(\cdot\mid x)$ over targets, from which a prediction is derived via a suitable inference (or decoding) procedure.
-Perhaps the most common inference strategy is to choose the mode of $\hat{p}(y\mid x)$:
-$\hat{y}(x):=\operatorname*{argmax}_{y\in Y}\hat{p}(y\mid x).$ (1)
-In practice, one may approximate the mode by employing greedy decoding or beam search, or sampling multiple candidates with a temperature and picking the among them the one with the highest likelihood score (Naseh et al., 2023).
+A language model (LM) takes a string $x$ as input and predicts an output $\hat{y}\in V^{*}$.
+Typically, the LM first produces a distribution $\hat{p}(\cdot\,|\,x)$ over targets.
+In a slight abuse of notation, we use $\hat{p}(y\,|\,x)\stackrel{{\scriptstyle\cdot}}{{=}}\hat{p}({\tt str}(y)\,|\,x)$ to denote the conditional probability of a numerical output $y$ given input $x$.
+Note that even for problems where numerical targets are expected, an LM may return a non-zero probability to non-numerical targets.
+A prediction from an LM is typically derived via a suitable inference (or decoding) procedure. Perhaps the most common inference strategy is to choose the mode of $\hat{p}(\cdot\,|\,x)$:
+$\hat{y}(x):=\operatorname*{argmax}_{y\in V^{*}}\hat{p}(y\,|\,x).$ (1)
+In practice, one may approximate the mode via greedy decoding or beam search, or sampling multiple candidates and picking the among them the one with the highest likelihood score (Naseh et al., 2023).
+In principle, the extracted target may not be numerical.
+In such cases, a possible strategy is to resort to predicting a default numerical value such as $0.0$.
+In practice, we find the targets from high-quality LLMs tend to be numerical even under zero-shot settings, and so converting most likely targets from $V^{*}$ to $Y$ is usually possible.
 The quality of an LM’s prediction is measured by some evaluation metric $m(y,\hat{y})$, where we assume that higher values are better.
 While the exact match (EM), given by $m(y,\hat{y})=\mathds{1}({y=\hat{y}})$, is a commonly used evaluation metric, there are a range of other metrics popularly used to evaluate LMs.
-These include the (negative) squared error $m(y,\hat{y})=-(y-\hat{y})^{2}$ or absolute error $m(y,\hat{y})=-|y-\hat{y}|$ in the case of numerical targets, or text-based metrics such as $F_{1}$ or BLEU.
-A natural goal is to then choose the inference strategy $\hat{y}(x)$ to maximize the metric $m$ of interest, i.e., to maximize the following expected utility:
+These include the (negative) squared error $m(y,\hat{y})=-(y-\hat{y})^{2}$ or absolute error $m(y,\hat{y})=-|y-\hat{y}|$ for regression tasks.
+A natural goal is to then choose the inference strategy $\hat{y}(x)$ to maximize the metric $m$ of interest, i.e., to maximize the expected utility:
 $\mathbb{E}_{(x,y)\sim D}\left[m(y,\hat{y}(x))\right].$ (2)
-As we shall see below, for many choices of metric $m(y,\hat{y}(x))$, picking the mode of the predicted distribution (1) can be sub-optimal for (2).
-Example 1.
-Consider the task of predicting the star rating (on the scale 11115555) associated with a review text.
-Suppose $m(y,\hat{y})$ is the negative absolute error between true and predicted star ratings.
-Given the review text ‘‘This keybord is suitable for fast typers’’, suppose the responses and the associated probabilities from an LM are {‘‘1’’: 0.3,  ‘‘2’’: 0.0,  ‘‘3’’: 0.3,  ‘‘4’’: 0.0,  ‘‘5’’: 0.4}.
+For many choices of metric $m(y,\hat{y}(x))$, picking the mode of the predicted distribution (1) can be sub-optimal for (2).
+As an example, consider predicting the star rating (on the scale 11115555) associated with a review text.
+Suppose $m(y,\hat{y})$ is the negative absolute error between the true and predicted ratings.
+Given the review text ‘‘This keybord is suitable for fast typers’’, suppose the LM responses and the associated probabilities are {‘‘1’’: 0.3,  ‘‘2’’: 0.0,  ‘‘3’’: 0.3,  ‘‘4’’: 0.0,  ‘‘5’’: 0.4}.
 The mode of the predicted probabilities is ‘‘5’’.
-In contrast, the maximizer of (2) is the median rating ‘‘3’’.
-We provide real examples for Amazon reviews of learn probability distributions and the label in Figure 2 (Appendix).
-Example 2.
-Consider a reading comprehension task, where we use the $F_{1}$ score as the evaluation metric $m(y,\hat{y})$, defined by the harmonic mean of $\text{recall}(y,\hat{y})=\frac{|y\cap\hat{y}|}{|y|}$ and $\text{precision}(y,\hat{y})=\frac{|y\cap\hat{y}|}{|\hat{y}|}$.
-Suppose for the question ‘‘What is the hottest month in the year’’, the responses and associated probability from an LM are {‘‘July’’: 0.25,  ‘‘July 2023’’: 0.23,  ‘‘Month of July’’: 0.24,  ‘‘May’’: 0.28}.
-The mode of this distribution is ‘‘May’’; whereas the maximizer of (2) is ‘‘July’’.
+In contrast, the maximizer of (2) is the median ‘‘3’’.
+(a) It is a nice color of black and my husband likes how it feels in his hand.
+(b) This item is a good idea. However, Unless the ear canal is reasonably deep (…) it’s of no use. The plastic hooks that come with it are hard and too small (…). Might be good for children.
+(c) One of the sides is made for apple products, the other is just standard usb. Both will work with apple products, just one side (the A side) charges faster. Other than that, it’s fantastic. :D
+Figure 2: Examples from the Amazon dataset and the corresponding: human annotations and samples from the model.
+We find that in many cases, taking into account the model distribution (i.e.
+a mean of the distribution) allows for a prediction closer to the annotation than simply taking the mode of the distribution.
+In Figure 2, we report examples from the Amazon dataset and the corresponding human annotations and samples from the model.
+Notice how samples cover significant proportions of the ratings.
+We find that the samples end up in the vicinity of the human annotation, and thus in many cases taking a mean over samples helps improve the prediction over the mode.
+Problem Label Pred. Metric Optimal rule $\Phi(\hat{y}_{1},\ldots,\hat{y}_{K},\hat{p}_{1},\ldots,\hat{p}_{K},\alpha)$
+Classification $[K]$ $[K]$ $\mathds{1}({y=\hat{y}})$ $\hat{y}(x):=\operatorname*{argmax}_{y}\,p(y\,|\,x)$ $\hat{y}_{i}$ s.t. $i=\operatorname*{argmax}_{j}\,\hat{p}_{j}$
+[…]
+Ordinal regression $[K]$ $[K]$ $-|y-\hat{y}|$ $\hat{y}(x):=\text{median}[p(\cdot\,|\,x)]$ $\hat{y}_{i}$ s.t. $i=\text{median}(\hat{p}^{\alpha}_{1},\ldots,\hat{p}^{\alpha}_{K})$
+Bi-partite ranking ±$\pm 1$ $\mathbb{R}$ AUC ($c_{y,y^{\prime}}=1$) $\hat{y}(x):=p(y=+1|x)$ $\sum_{i}\frac{\hat{p}_{i}^{\alpha}\cdot\mathds{1}(\hat{y}_{i}=1)}{\sum_{j}\hat{p}_{j}^{\alpha}\cdot\mathds{1}(\hat{y}_{j}=1)}$
+[…]
+Table 1: Optimal decision rule for varying: label space, model prediction space and evaluation metric.
+We denote $[K]=\{1,\ldots,K\}$.
+The final column shows the empirical rule as a function of sampled outputs, corresponding scores, and a rescaling temperature $\alpha$ (see Section 3.3).
 3 Metric-aware LLM inference
-We seek to design decoding strategies that maximize the expected utility in (2).
-3.1 Bayes-optimal inference
-In an idealized setting, where we have access to the true conditional probabilities $p(\cdot\mid x)$, it is straight-forward to show that the maximizer of (2) is:
-$\hat{y}^{*}(x)\in\operatorname*{argmax}_{y^{\prime}\in Y}\,\mathbb{E}_{y\sim p(\cdot\mid x)}\left[m(y,y^{\prime})\right].$ (3)
-When $m$ is the EM metric, the optimal inference strategy is $\hat{y}^{*}(x)\in\operatorname*{argmax}_{y\in Y}\,p(y\mid x),$ which is what common approaches such as greedy decoding seek to approximate.
-In general, however, the optimal decoding strategy can have a very different form.
-For example, for evaluation metrics over numerical targets such as the squared error or the absolute error, the optimal inference strategy is to simply take the mean or median of $p(\cdot|x)$ (Bishop, 2006).
-Table 1 summarizes these optimal rules for different metrics.
-3.2 Approximating the optimal solution
+3.1 Minimum Bayes risk decoding
+We seek to design decoding strategies that maximize the expected utility in (2). Ideally, if we had access to the true conditional probabilities $p(\cdot\,|\,x)$, the maximizer of (2) is given by:
+$\hat{y}^{*}(x)\in\operatorname*{argmax}_{y^{\prime}\in Y}\,\mathbb{E}_{y\sim p(\cdot\,|\,x)}\left[m(y,y^{\prime})\right].$ (3)
+When $m$ is the EM metric, the optimal inference strategy is $\hat{y}^{*}(x)\in\operatorname*{argmax}_{y\in Y}\,p(y\,|\,x),$ which is what common approaches such as greedy decoding seek to approximate.
+In general, however, the optimal decoding strategy can have a very different form, and the mode of $p(\cdot|x)$ has been shown to be suboptimal on generation tasks (Eikema and Aziz, 2020).
+For example, as shown in Table 1, for evaluation metrics over numerical targets such as the squared error or the absolute error, the optimal inference strategy is to take the mean or median of $p(\cdot|x)$ (Bishop, 2006).
+3.2 Closed-form optimal solution
 In practice, we mimic the Bayes-optimal solution in (3) with two approximations.
-First, we replace the true conditional distribution ${p}(\cdot\mid x)$ with the LM’s predicted distribution $\hat{p}(\cdot\mid x)$.
-This is a reasonable approximation when the LM is pre-trained with next-token prediction task and softmax cross-entropy loss: the latter is a strictly proper loss, whose minimizer under an unrestricted hypothesis class is the true conditional distribution $p(y\mid x)$ (Gneiting and Raftery, 2007).
-Second, we estimate the expectation in (3) by sampling $K$ outputs from $\hat{p}(\cdot\mid x)$ with a sampling temperature $T$, and then computing:
+First, we replace the true conditional distribution ${p}(\cdot\,|\,x)$ with the LM’s predicted distribution $\hat{p}(\cdot\,|\,x)$.
+This is a reasonable approximation when the LM is pre-trained with next-token prediction objective based on the softmax cross-entropy loss; the latter is a strictly proper loss, whose minimizer under an unrestricted hypothesis class is the true conditional distribution $p(y\,|\,x)$ (Gneiting and Raftery, 2007).
+Second, we estimate the expectation in (3) by sampling $K$ outputs from $\hat{p}(\cdot\,|\,x)$, and then computing:
 $\hat{y}(x)\in\operatorname*{argmax}_{y^{\prime}\in Y}\,\sum_{i=1}^{K}m(y_{i},y^{\prime}).$ (4)
-As noted in Table 1, the above maximization can be computed in closed-form for some common metrics.
-In applications where we neither have a closed-form solution, nor can enumerate $S$ feasibly, we may solve (4) approximately by maximizing over a fixed subset $C\subset Y$ of candidate outputs:
-$\hat{y}(x)\in\operatorname*{argmax}_{y^{\prime}\in C}\,\sum_{i=1}^{K}m(y_{i},y^{\prime}).$ (5)
-The choice of $C$ would depend on the task and metric at hand.
-For example, one simple choice is for $C$ to be the $K$ sampled outputs, i.e., $C=\{y_{1},\ldots,y_{K}\}$.
-In Appendix A.1, we described how we choose $C$ for text generation tasks.
-The quality of our inference strategy depends on the sampling temperature $T$, the number of samples $K$, and the choice of $C$.
-Note that sampling can be done efficiently by caching the Transformer activations for the input prefix when generating different targets; the Transformer computation needs to be repeated only for the different sampled targets.
+Even with these approximations, maximizing (4) over all outputs $Y$ is intractable in general.
+Prior literature on MBR for metrics like BLEU heuristically perform maximization over a small set of candidates (Ehling et al., 2007; Bertsch et al., 2023).
+In this paper, we consider regression and scoring metrics, for which the above maximization can be computed in closed-form.
+As shown in Table 1, these solutions can be estimated by computing simple statistics from the sampled responses, such as the sample mean $\hat{y}(x)=\frac{1}{K}\sum_{i=1}^{K}y_{i}$ for the squared error.
+We refer to this approach as Regression-aware Inference with LLMs (RAIL).
 3.3 Post-hoc temperature scaling
-As mentioned, when sampling from $\hat{p}(\cdot\mid x)$, it often helps to apply a temperature scaling to the LM logits to control the diversity of the sampled outputs.
-This is particularly important in our procedure where we wish to approximate expectations over $\hat{p}(\cdot|x)$ using a few samples.
-In practice, one may sample from $\hat{p}(\cdot\mid x)$ with temperature $T=1$, and apply temperature scaling in a post-hoc manner by employing a weighted version of the objective in (4):
-$\hat{y}(x)\in\operatorname*{argmax}_{y^{\prime}\in Y}\,\sum_{i=1}^{K}\left(\hat{p}(y_{i}\mid x)\right)^{\alpha}\cdot m(y_{i},y^{\prime}),$ (6)
+When sampling from $\hat{p}(\cdot\,|\,x)$, it often helps to apply temperature scaling to the LM logits to control the sampled outputs’ diversity.
+This is particularly important in our procedure, where we wish to approximate expectations over $\hat{p}(\cdot|x)$ using a few samples.
+In practice, one may sample from $\hat{p}(\cdot\,|\,x)$ with temperature $T=1$, and apply temperature scaling in a post-hoc manner by employing a weighted version of the objective in (4):
+$\hat{y}(x)\in\operatorname*{argmax}_{y^{\prime}\in Y}\,\sum_{i=1}^{K}\left(\hat{p}(y_{i}|x)\right)^{\alpha}\cdot m(y_{i},y^{\prime}),$ (5)
 where $\alpha$ can be seen as the temperature scaling parameter.
-The above summation is a (scaled) estimate of $\mathbb{E}_{y\sim\hat{p}(\cdot\mid x)}\left[\hat{p}(y\mid x)^{\alpha}\cdot m(y,y^{\prime})\right]$.
-For probabilities $\hat{p}(y_{i}\mid x)\propto\exp(f(x,y_{i}))$ defined by logits $f(x,y_{i})$, this is equivalent to computing the expectation under the temperature-scaled distribution $\hat{p}_{\alpha}(y\mid x)\propto\exp((1+\alpha)\cdot f(x,y))$, albeit a normalization factor.
-We consider an analogous weighting scheme for the plug-in estimators of the closed-form solutions reported in Table 1.
-Problem Label space $Y$ Metric Optimal decision rule
-Classification $1,\ldots,K$ $\mathds{1}({y=\hat{y}})$ $\hat{y}(x):=\operatorname*{argmax}_{y}[p(y\mid x)]$
-Regression $\mathbb{R}$ $-(y-\hat{y})^{2}$ $\hat{y}(x):=\mathbb{E}_{y\sim p(\cdot\mid x)}[y]$
-Ordinal regression $1,\ldots,K$ $-|y-\hat{y}|$ $\hat{y}(x):=\text{median}[p(\cdot\mid x)]$
-Question answering $V^{*}$ $F_{1}(y,\hat{y})$ $\hat{y}(x):=\operatorname*{argmax}_{y^{\prime}}\,\mathbb{E}_{y\sim p(\cdot\mid x)}\left[m(y,y^{\prime})\right]$
-Bi-partite ranking ±$\pm 1$ AUC $\hat{y}(x):=p(y=+1|x)$
-Multi-partite ranking $1,\ldots,K$ $\text{AUC}^{\text{multi}}$  with costs $c_{y,y^{\prime}}=|y-y^{\prime}|$ $\hat{y}(x):=\mathbb{E}_{y\sim p(\cdot\mid x)}[y]$
-Table 1: Optimal inference for different evaluation metrics.
-For the first three, we have a closed-form solution.
-For $F_{1}$, we approximate the optimal solution by optimizing over a subset of candidates $C\subset Y$3.2).
-See (§3.5) for definitions of AUC.
-For the problems in rows 1–4, the LLM predicts an output in label space $Y$; for the ranking problems in rows 5–6, the LLM predicts a real-valued score, which is then used to rank the samples.
-model size greedy decode MALI
-argmax mean $*{}^{*}$
-STSB XXS 1.078 1.448 1.007 1.028
-S 0.685 1.019 0.641 0.649
-L 0.628 0.989 0.580 0.610
-argmax median $*{}^{*}$
-Amazon reviews XXS 0.495 0.826 0.493 0.474
-S 0.301 0.444 0.299 0.285
-L 0.294 0.541 0.298 0.291
-argmax $F_{1}$ ${F_{1}}^{*}$
-Trivia-QA XXS 0.314 0.178 0.307 0.302
-S 0.620 0.636 0.650 0.678
-L 0.886 0.887 0.880 0.886
-Table 2: Root mean squared error (RMSE) on STSB dataset (lower is better), Mean absolute error (MAE) on Amazon reviews dataset (lower is better), and $F_{1}$ metrics on Trivia-QA dataset (higher is better) from PaLM-2 models of varying size.
-We sample $16$ targets for the argmax and MALI methods.
-All methods rely on $T=1$ temperature for sampling, and the weighted approaches correspond to an effective temperature $T=\frac{1}{4}$.
-See Tables 5, 6, 7 in the Appendix for results over more temperatures and evaluation metrics (cost weighted multi-class AUC and Pearson correlation).
-3.4 Efficiency
-Note that sampling can be done efficiently by caching the Transformer activations for the input prefix when generating different targets.
-The Transformer computation needs to be repeatedly done only for the different sampled targets.
-In practice, when the prefix is long compared to the generated targets (e.g.
-for scoring tasks, where the target score can be just a few tokens length, whereas the prefix can be long as it contains the input text), a forward pass for the prefix tends to take most of the compute time.
-Moreover, we generate multiple samples simultaneously, and so, we do not incur a higher cost from generating multiple targets.
-3.5 Ranking metrics
-Apart from the settings where evaluation metrics are decomposable, i.e., evaluation metrics are a result of aggregation of per-example results from evaluation, there are also metrics which are non-decomposable.
-An example for such a metric are ranking metrics, where the goal is to correctly order examples, as opposed to predict correct scores or classes.
-Ranking metrics are appropriate for many important tasks, including ads click-through rate prediction or recommender systems.
-Below we review different types of ranking metrics and the corresponding optimal decision rules.
-Bi-partite ranking
-In bi-partite ranking, the goal is to order examples from two classes (and so the label set $Y={\pm 1}$) in such a way that examples from the positive class are above the examples from the negative class.
-In this case The LLM predicts real valued scores, and so the target space need not be the same as the label space.
-One common metric used for evaluating bi-partite ranking is the area under the ROC curve (AUC), which is proportional to the number of wrongly ordered pairs of examples in the ranking (Cortes and Mohri, 2003):
-$\text{AUC}(\hat{y})\,=\,1-\mathbb{E}\Big{[}\mathds{1}(\hat{y}(x)<\hat{y}(x^{\prime}))\,+\,\frac{1}{2}\cdot\mathds{1}(\hat{y}(x)=\hat{y}(x^{\prime})\,\Big{|}\,y=+1,y^{\prime}=-1\Big{]}$
-It can be shown that the optimal decision rule for AUC takes the form of (any strictly monotone transformation of) $\hat{y}^{*}(x)=p(y=+1|x)$ (Clémençon et al., 2008; Uematsu and Lee, 2015).
-Multi-partite ranking
-Multi-partite ranking is an extension to the bi-partite ranking, where the ranking is over examples from multiple categories.
-Several multi-category extensions of binary AUC have been considered, for example by counting wrongly ordered pairs of examples in the ranking according to the category indices and weighing them according to predetermined costs (Uematsu and Lee, 2015):
-$\text{AUC}^{\text{multi}}(\hat{y})\,=\,1-\mathbb{E}\Big{[}c_{y,y^{\prime}}\cdot\Big{(}\mathds{1}(\hat{y}(x)<\hat{y}(x^{\prime}))\,+\,\frac{1}{2}\cdot\mathds{1}(\hat{y}(x)=\hat{y}(x^{\prime})\Big{)}\,\Big{|}\,y>y^{\prime}\Big{]}$
-Uematsu and Lee show that when the costs are equal to the difference between category indices, i.e., $c_{y,y^{\prime}}=|y-y^{\prime}|$, an optimal solution is to use mean distribution over scores: $\hat{y}^{*}(x)=\mathbb{E}_{y\sim p(\cdot|x)}\left[y\right]$.
-We report results on the cost weighted multi-class AUC in Table 6 in Appendix.
+The above summation is a (scaled) estimate of $\mathbb{E}_{y\sim\hat{p}(\cdot\,|\,x)}\left[\hat{p}(y\,|\,x)^{\alpha}\cdot m(y,y^{\prime})\right]$.
+For probabilities $\hat{p}(y_{i}\,|\,x)\propto\exp(f(x,y_{i}))$ defined by logits $f(x,y_{i})$, this is equivalent to computing the expectation under the temperature-scaled distribution $\hat{p}_{\alpha}(y\,|\,x)\propto\exp((1+\alpha)\cdot f(x,y))$, modulo a normalization factor.
+We consider an analogous weighting scheme for the plug-in estimators of the closed-form solutions in Table 1.
+Algorithm 1 outlines the RAIL procedure, with both a sampling temperature $T$ and an effective temperature $T^{\prime}$ as inputs.
+The algorithm first draws samples from an LLM with the sampling temperature $T$; next, to arrive at an effective temperature $T^{\prime}$, it performs post-hoc scaling by a factor $\alpha=\frac{T}{T^{\prime}}-1$.
+In principle, temperature scaling may not be necessary if $\hat{p}(y\,|\,x)$ accurately estimated the true probability $p(\cdot\,|\,x)$.
+However, in practice due to imperfect approximation and finite sample size errors, we find it useful to employ.
+Indeed, temperature scaling has also been found to be beneficial in prior MBR works (Yan et al., 2022).
+3.4 Extension to multi-partite ranking
+Our metric-aware decoding proposal also applies to scoring tasks, where the label space $Y$ is discrete, e.g.
+$\{1,\ldots,K\}$, but we require the LLM to predict real-valued scores $\hat{y}(x)\in\mathbb{R}$ for each prompt $x$ such that prompts with higher labels receive a higher score.
+One typically measures the performance of $\hat{y}(x)$ using a pairwise ranking metric such as the multi-partite area under the ROC curve (AUC-ROC) (Uematsu and Lee, 2015):
+[…]
+which penalizes the scorer $\hat{y}$ by $c_{y,y^{\prime}}$ whenever it mis-ranks a pair $(x,x^{\prime})$ with $y>y^{\prime}$. In experiments, we refer to AUC-ROC as AUC for brevity.
+Despite AUC-ROC being non-decomposable (not a summation of per-example results), Uematsu and Lee (2015)[Corollary 1] show that when the costs are the difference between the labels, i.e., $c_{y,y^{\prime}}=|y-y^{\prime}|$, the optimal scorer admits a closed-form solution given by the expected label under distribution $p(\cdot|x)$$\hat{y}^{*}(x)=\mathbb{E}_{y\sim p(\cdot|x)}\left[y\right]$.
+One can thus readily apply our RAIL approach to estimate this solution from sampled responses.
+Moreover, from the Neyman-Pearson lemma, the same optimal solution applies to the AUC-PR evaluation metric Clémençon and Vayatis (2009).
 4 Experiments and Discussion
-We compare our metric-aware inference approach against standard inference strategies on NLP tasks with different evaluation metrics.
-Datasets. We summarize the datasets below, and list the prompts used in Table 4 (Appendix).
-  1. (i)
-    STSB: The Semantic Textual Similarity Benchmark (Cer et al., 2017) comprises of sentence pairs human-annotated with a similarity score from 0 to 5.
-As it is a regression task, we consider evaluation with root mean squared error.
-  2. (ii)
-    US Amazon reviews: We evaluate on a sample of $1500$ examples from the US Amazon reviews dataset.
-From the review text we aim to predict the 5-star rating corresponding to a product review (Ni et al., 2019).
-Since the task is in the form of ordinal regression, we use mean absolute error as the metric (Fathony et al., 2017).
-  3. (iii)
-    TriviaQA is a reading comprehension dataset comprising (question, answer, reference text) triplets, where a question is answered against a refence text (Joshi et al., 2017).
-The evaluation is the $F_{1}$ score (see Example 2).
-We use a sample of $1500$ examples for evaluation.
+We experimentally evaluate our proposed approach on NLP tasks with different evaluation metrics.
+Datasets.
+We use two datasets: (i) Semantic Textual Similarity Benchmark (STSB) (Cer et al., 2017), which comprises of sentence pairs human-annotated with a similarity score from 0 to 5; since this is a regression task, we evaluate with the root mean squared error.
+(ii) US Amazon reviews, where we aim to predict the 5-star rating for a product review (Ni et al., 2019); since the task is in the form of ordinal regression, we use mean absolute error as the evaluation metric (Fathony et al., 2017).
+We list the prompts used in Table 6 (Appendix).
+In each case, we evaluate on samples of $1500$ examples.
+model size greedy decode RAIL
+argmax mean
+[…]
+S 0.685 1.019 0.649
+L 0.628 0.989 0.610
+argmax mean
+[…]
+S 0.895 0.820 0.953
+L 0.905 0.827 0.961
+argmax median
+[…]
+S 0.301 0.444 0.285
+L 0.294 0.541 0.291
+Table 2: Comparison of inference strategies on PaLM-2 models for different datasets and metrics.
+We draw $16$ samples with an effective temperature of $T=\frac{1}{4}$ (via post-hoc scaling).
+In Appendix D, we report results for variants of MALI with no post-hoc scaling (Table 4), and results of tuning the temperature using a held-out set, along with confidence intervals (Table 11).
 Models.
-We consider two instruction-tuned model families: FLAN-T5 (Chung et al., 2022) and PaLM-2 (Google and et al., 2023).
+We consider two instruction-tuned model families: PaLM-2 (Anil et al., 2023) and FLAN-T5 (Chung et al., 2022).
 We report results across different model sizes and temperatures.
@@
 Methods.
-We evaluate the following methods (i) greedy decoding, (ii) sampling $K$ candidates and picking the one with the maximum likelihood (argmax), (iii) the proposed MALI approach on the same $K$ samples, and (iv) the temperature scaled variant of MALI in §3.3 (denoted by a ‘*’).
-For (iv), we choose $\alpha$ so that the effective temperature is $\frac{1}{4}$.
-For the tasks involving the squared and absolute error metrics, we use the closed-form decision rules in Table 1.
-For the $F_{1}$ metric, we solve (5) over a candidate set $C$, which we choose to contain the $K$ samples and additional targets derived from them (details in Appendix A.2).
+We evaluate the following methods: (i) greedy decoding, (ii) a baseline inspired from the self-consistency decoding of sampling $K$ candidates and picking the one with the maximum likelihood (argmax) (Wang et al., 2023a), (iii) the proposed RAIL approach on the same $K$ samples, and (iv) RAIL with temperature scaling (§3.3).
+For (iv), we choose $\alpha$ yielding effective temperature $\frac{1}{4}$.
 Metric-aware inference helps.
 In Table 2, we report results across datasets and model sizes.
-We notice that MALI improves over baselines across all model sizes on STSB and Amazon reviews (for model size S, we see that median performs very similarly to the most likely generated sample).
-On Trivia QA, we find that on XXS and S, optimizing for $F_{1}$ helps, and for L, simply using the most likely generated sample works most effectively.
-We see diminishing improvements with larger models for Trivia QA and Amazon datasets.
-This coincides with a lowering entropy in predictions with increasing model size for Trivia QA and Amazon datasets, whereas STSB does not exhibit this phenomenon as strongly: see Table 8 (Appendix) for empirical entropy estimates.
-We also evaluate MALI with varying number of samples $K$.
-As seen in Table 9 (Appendix), on STSB with temperature $\frac{1}{4}$, even with as few as two samples, MALI starts to show improvements over greedy decoding.
-Sampling versus scoring.
-So far, when estimating the prediction maximizing the (2), we have used sampling from the LM distribution (see §3.2).
-Alternatively, if the targets are from a narrow interval (e.g., on the STSB dataset, the values are in the interval $[0,5]$), one can find the probabilities for targets at fixed intervals (e.g.
-$0,0.5,1.0,\ldots,5.0$), and thus find the solutions from Table 1 and more generally from (3).
-In Table 3, we report results from FLAN-T5 on the STSB dataset for both scoring and sampling based estimates for MALI, where the scoring based estimator is based on $11$ equally spaced targets.
-We find that both sampling and scoring lead to MALI improving over choosing the most likely prediction.
-Further, we note that sampling is a more effective strategy than scoring the equally spaced targets.
-model greedy scoring sampling
-FLAN-T5 S 4.419 2.407 2.275
-FLAN-T5 L 0.455 0.410 0.373
-FLAN-T5 XL 0.508 0.549 0.457
-Table 3: Squared error (SE) on STSB from FLAN-T5 models of varying size. Temperature $0.5$ was used for generating targets for the sampling approach.
-Conclusions In this work we show how utilizing the output distribution modeled by LLMs in the form of our MALI methods can bring improvements across NLP tasks.
-Our work also points at importance of further understanding and improving of the quality of model output distribution and calibration, which we found crucial for the MALI methods to work well.
+We notice that RAIL consistently improves over baselines.
+To better measure the sensitivity of the results to the choice of temperature, we report additional results in Table 11 in Appendix D, where we use a held-out validation set to tune the temperature, and find the trends to be consistent.
+Sampling versus enumeration.
+So far, when estimating the maximizer to equation (2), we have used sampling from the LM distribution (see §3.2).
+Alternatively, if the targets are from a narrow interval (e.g., on STSB, the values are in the interval $[0,5]$), one can score the model for targets enumerated at fixed intervals (e.g.
+$0,0.5,1.0,\ldots,5.0$), and compute estimates for solutions in Table 1.
+In Table 3, we report results from FLAN-T5 on the STSB dataset for RAIL with both sampling and enumeration based estimates, where the latter is based on $11$ equally spaced targets.
+Both sampling and enumeration lead to RAIL improving over choosing the most likely target, with sampling having an edge.
+The reason sampling performs better than enumeration may be that sampling is able to better explore the high density regions of the output probability space, as we detail in Appendix E.
+model greedy enumeration sampling
+FLAN-T5 S 2.102 1.551 1.508
+FLAN-T5 L 0.675 0.640 0.611
+FLAN-T5 XL 0.713 0.741 0.676
+Table 3: RMSE on STSB with FLAN-T5 across RAIL variants (enumeration vs sampling). The sampling approach uses a temperature of $0.5$.
+Role of model size.
+We find that the benefit from our technique reduces as the models increase in size.
+This sometimes coincides with a lowering entropy in predictions with increasing model size (see, e.g., results on Amazon in Table 7 in Appendix).
+We note this is consistent with prior works on MBR, which observed that as the model gets better, the optimal decision rule for EM (approximated by greedy decoding) performs comparable to the that for other metrics (Schluter et al., 2012).
+We stress that the gains we get with small and medium-sized models are still of large practical importance, especially in applications where deploying very large models is prohibitively expensive.
+5 Conclusions
+In this work we show how adopting MBR to regression and scoring tasks, and thus utilizing the output distribution modeled by LLMs in the form of our RAIL methods can bring improvements.
+Our work also points at importance of further understanding and improving of the quality of model output distribution and calibration, which we found crucial for the RAIL methods to work well.
+In the future, we wish to extend our approach to other less-explored evaluation metrics in the MBR literature.
 6 Limitations
 There are multiple limitations of our work.
-First, we evaluate our proposed methods on multiple text datasets with numerical and text targets, however, many more types of outputs can be considered, including time series targets, long text targets etc.
-Next, it would be interesting to more systematically analyze how to solve the objective from ((6)) over many samples for text outputs for metrics like $F_{1}$ or BLEU, e.g.
+First, we evaluate our proposed methods on multiple text datasets with numerical and text targets, however, many more types of outputs can be considered, including the time series targets.
+Next, it would be interesting to more systematically analyze how to efficiently solve the objective from (5) over many samples for text outputs for metrics like $F_{1}$ or BLEU, e.g.
 by means of dynamic programming.
@@
 No additional user data was collected or released as part of this work.
-All models used are publicly available and already pretrained, and no finetuning was conducted for any experiments.
+All models used are publicly available and already pretrained, and no fine-tuning was conducted for any experiments.
 Instead, all experiments relied on running inference experiments with the models over several thousands of examples.
@@
 8 Acknowledgements
-We are thankful to Ziwei Ji for the helpful feedback.
-Appendix A Appendix
-A.1 Candidate set construction
-One simple choice for the candidate set $C$ could be take the $K$ sampled outputs, i.e., $C=\{y_{1},\ldots,y_{K}\}$.
-One may additionally include in this set transformations on each $y_{i}$ $y_{i}$ or new candidates formed from combining two or more of the samples.
-For example, in applications such as reading comprehension or question-answering, where the output is list of keywords that constitute an answer to a question, one often uses a metric such as F1 that balances between precision and recall.
-In such cases, after sampling, one may additionally include samples formed by concatenating pairs of sampled outputs, i.e., ${\tt concat}(y_{i},{\tt delim},y_{j}),\forall i\neq j$.
-These concatenated answers have the effect of increasing recall, at the cost of lower precision.
-We follow that procedure for the Trivia-QA experiments in §4.
-A.2 Additional details
-In Table 4 we report the prompts we used in our experiments for zero-shot inference.
-For all datasets, we use validation splits, and where not available, we use the first $1500$ examples from the train split.
-Dataset Prompt
-STSB
-What is the sentence similarity between the following two sentences measured on a scale of 0 to 5: {Sentence #1}, {Sentence #2}.
-The similarity measured on a scale of 0 to 5 with 0 being unrelated and 5 being related is equal to
-Amazon reviews
-What is the rating corresponding to the following review in the scale of 1 to 5, where 1 means negative, and 5 means positive? Only give a number from 1 to 5 with no text. Review: {Review} Rating:
-Trivia-QA
-Answer the following question without any additional text. Question: {Question}. Answer:
-Table 4: Prompts used for different datasets. Curly braces denote inputs specific to an input example.
-A.3 Additional experiments
-In Table 8 we report empirical entropy estimates as measured based on the $16$ samples generated from the model.
-We find that entropy decreases as model size increases.
-We observe a particularly sharp decrease in entropy for the Amazon reviews and Trivia-QA datasets, where for larger model sizes we don’t find improvements from MALI approaches.
-In Table 5 we report RMSE on STSB dataset, MAE on Amazon reviews dataset, and $F_{1}$ metrics on Trivia-QA dataset from PaLM-2 models of varying size across multiple temperature values.
-We find improvements over baselines on STSB and Amazon reviews datasets for most temperatures.
-For Trivia-QA, we find improvements for XXS and S models for some temperatures, and for L, we don’t find a difference from our methods due to low entropy in the responses (see Table 8).
-In Table 7 we additionally report Pearson correlation metrics on STSB, confirming the results of MALI improving over autoregressive inference.
-Lastly, in Table 6 we report cost weighted multi-class AUC with costs corresponding to the difference between the annotated labels: $|y_{1}-y_{2}|$.
-We find on both STSB and Amazon reviews datasets that the optimal decision rule (mean over the distribution) improves over the baselines.
-In Table 9, we report the impact of the number of samples on the results.
-We note that there is an improvement in the results with the increase in the number of samples, however beyond 8 samples there is a diminishing improvement in practice.
-In Figure 2 we report examples from the Amazon dataset and the corresponding: human annotations and samples from the model.
-Notice how samples cover significant proportions of the ratings.
-We find that the samples end up in the vicinity of the human annotation, and thus in many cases taking a mean over samples helps improve the prediction over the mode.
-(a) It is a nice color of black and my husband likes how it feels in his hand.
-(b) This item is a good idea. However, Unless the ear canal is reasonably deep (…) it’s of no use. The plastic hooks that come with it are hard and too small (…). Might be good for children.
-(c) One of the sides is made for apple products, the other is just standard usb. Both will work with apple products, just one side (the A side) charges faster. Other than that, it’s fantastic. :D
-Figure 2: Examples from the Amazon dataset and the corresponding: human annotations and samples from the model.
-We find that in many cases, taking mean over model samples is closer to the annotation than taking the mode of the distribution.
+We are thankful to Changsheng Jiang for experiments and analyses of early versions of our technique. We are also thankful to Ziwei Ji and Shankar Kumar for helpful feedback and comments.
+Appendix A Further related work
+Minimum Bayes risk decoding.
+As noted in the introduction, prior work on MBR has considered optimizing for common metrics in the machine translation and text generation literature.
+The closest to our paper is the work of Wang et al.
+[2023a], who considered sampling from the model distribution using chain-of-thought (CoT) prompting, and showed how majority vote improves over the baseline on arithmetic and reasoning tasks.
+Other works explored different aspects of MBR, including: the role of the sampling algorithms [Freitag et al., 2023, Cheng and Vlachos, 2023], the interaction with label smoothing [Yan et al., 2022], and how it generalizes other techniques [Suzgun et al., 2022, Bertsch et al., 2023].
+Finkelstein and Freitag [2024] recently considered distillation of MBR solution to a student model, so as to avoid the overhead induced by MBR at inference time.
+A recent work also applied LLMs to time series forecasting, and constructed the final predictions by computing quantiles (e.g., median) over the samples [Gruver et al., 2023].
+One of the evaluation metrics for time series forecasting is the mean absolute error, for which the median can be shown to be a Bayes optimal decision rule (see Table 1).
+Fine-tuning for target task alignment.
+Previous works have considered approaches for aligning the models for target datasets.
+This includes fine-tuning of soft prompts on target datasets without losing generalization to other tasks [Wang et al., 2023b], and general fine-tuning on carefully tailored datasets for improved model robustness [Li et al., 2023].
+In our work, we focus on zero-shot setting where no fine-tuning is conducted.
+Fine-tuning for numerical tasks.
+Autoregressive fine-tuning of LLMs on numerical tasks with CoT has been found effective [Liu and Low, 2023].
+One line of work for modeling predictive tasks with pre-trained Transformer-based models is to add a regression head on top of the transformed/pooled encoded input tokens and fine-tune the resulting model on numerical targets using a regression loss.
+This is an approach which has been employed for encoder-based models (e.g., BERT), and has also been applied to encoder-decoder (e.g., T5) models [Liu et al., 2022], and these approaches could be extended to decoder models too.
+In a similar work, an embedding was extracted from a decoder model fine-tuned on modified attention mask and additional tasks [BehnamGhader et al., 2024].
+In this paper, we focus on the zero shot approaches, and we leave training approaches for future work.
 model size greedy decode T=$0.25$ T=$0.5$ T=$1.0$
@@
 For the weighted approaches, we fix the sampling temperature to $T=1$ and accordingly vary the $\alpha$ in (5) so as to arrive at the effective temperature equal to the value reported.
-model size greedy decode T=$0.25$ T=$0.5$ T=$1.0$
-argmax mean argmax mean argmax mean
-STSB XXS $0.797$ $0.755$ $0.797$ $0.755$ $0.797$ $0.755$ $0.797$
-XS $0.895$ $0.870$ $0.895$ $0.870$ $0.895$ $0.870$ $0.895$
-L $0.905$ $0.885$ $0.905$ $0.885$ $0.905$ $0.885$ $0.905$
-Amazon reviews XXS $0.87$ $0.894$ $0.925$ $0.866$ $0.94$ $0.788$ $0.942$
-XS $0.9$ $0.91$ $0.925$ $0.914$ $0.941$ $0.9$ $0.958$
-L $0.925$ $0.922$ $0.951$ $0.906$ $0.962$ $0.837$ $0.964$
-Table 6: Cost-weighted multi-partite AUC metrics on STSB and Amazon datasets (the higher the better).
-MALI methods improve over the baselines.
-See §3.5 for the definition of AUC we use.
-We assume costs to correspond to the difference between the annotated labels: $|y_{1}-y_{2}|$.
-model greedy decode T=$0.25$ T=$0.5$ T=$1.0$
-argmax mean argmax mean argmax mean
-PaLM-2 XXS $0.767$ 0.738 $\mathbf{0.790}$ 0.670 $\mathbf{0.790}$ 0.544 $0.786$
-PaLM-2 XS $0.898$ 0.878 $\mathbf{0.915}$ 0.852 $0.913$ 0.821 $0.910$
-PaLM-2 L $0.909$ 0.893 $0.920$ 0.881 $0.922$ 0.860 $\mathbf{0.923}$
-Table 7: Pearson correlation metrics on STSB. MALI methods improve over the baselines.
+model w/ pairs w/o pairs
+PaLM-2 XXS $0.302$ $0.295$
+PaLM-2 XS $0.678$ $0.670$
+PaLM-2 L $0.886$ $0.887$
+Table 5: Performance of RAIL (as evaluated by $F_{1}$) on TriviaQA with and without the inclusion of concatenated pairs in the candidate set.
+Appendix B Additional results on $F_{1}$ maximization on Trivia QA
+We extend our approach to the $F_{1}$ score evaluation metric.
+Consider a reading comprehension task, where the $F_{1}$ score is the evaluation metric $m(y,\hat{y})$, defined by the harmonic mean of $\text{recall}(y,\hat{y})=\frac{|y\cap\hat{y}|}{|y|}$ and $\text{precision}(y,\hat{y})=\frac{|y\cap\hat{y}|}{|\hat{y}|}$.
+To illustrate the task, suppose for the question ‘‘What is the hottest month in the year’’, the responses and associated probability from an LM are {‘‘July’’: 0.25,  ‘‘July 2023’’: 0.23,  ‘‘Month of July’’: 0.24,  ‘‘May’’: 0.28}.
+The mode of this distribution is ‘‘May’’; whereas the maximizer of (2) is ‘‘July’’.
+To optimize the $F_{1}$ metric, we solve (7) over a candidate set $C$, which we choose to contain the $K$ samples and additional targets derived from them.
+$\hat{y}(x)\in\operatorname*{argmax}_{y^{\prime}\in C}\,\sum_{i=1}^{K}m(y_{i},y^{\prime}).$ (7)
+While the $F_{1}$ score does not admit a closed-form solution, as is the case for the metrics listed in Table 1, we make an observation that its formulation allows for introducing a different form of efficiency.
+In particular, we notice that due to the trade-off between precision and recall in the $F_{1}$ score formulation, the following candidate set construction can lead to increasing recall at the expense of precision, thus providing a way to cheaply enumerate additional reasonable candidates.
+Dataset Prompt
+STSB What is the sentence similarity between the following two sentences measured on a scale of 0 to 5: {Sentence #1}, {Sentence #2}.
+The similarity measured on a scale of 0 to 5 with 0 being unrelated and 5 being related is equal to
+Amazon reviews What is the rating corresponding to the following review in the scale of 1 to 5, where 1 means negative, and 5 means positive?
+Only give a number from 1 to 5 with no text.
+Review: {Review} Rating:
+Table 6: Prompts used for different datasets. Curly braces denote inputs specific to an input example.
+Candidate set construction.
+One simple choice for the candidate set $C$ could be take the $K$ sampled outputs, i.e., $C=\{y_{1},\ldots,y_{K}\}$.
+One may additionally include in this set transformations on each $y_{i}$ or new candidates formed from combining two or more of the samples.
+For reading comprehension or question-answering applications, where the output is a list of keywords that constitute an answer to a question, one may additionally include samples formed by concatenating pairs of sampled outputs, i.e., ${\tt concat}(y_{i},{\tt delim},y_{j}),\forall i\neq j$.
+These concatenated answers have the effect of increasing recall, at the cost of lower precision.
+We follow that procedure for the Trivia-QA experiments.
+In Table 4, we provide results on Trivia-QA reading comprehension task [Joshi et al., 2017] with the proposed $F_{1}$-aware inference strategy.
+To additionally analyze the effectiveness of the candidate set augmentation, in Table 5 we compare the performance of RAIL (specifically the temperature scaled variant) with and without the inclusion of concatenated pairs in the candidate set.
+For both the XXS and S models, the inclusion of concatenated pairs is seen to yield a significant improvement in $F_{1}$-score.
+Appendix C Additional details
+In Table 6 we report the prompts we used in our experiments for zero-shot inference.
+For all datasets, we use validation splits, and where not available, we use the first $1500$ examples from the train split.
+Appendix D Additional experiments
+In Table 7 we report empirical entropy estimates as measured based on the $16$ samples generated from the model.
+We find that entropy decreases as model size increases.
+We observe a particularly sharp decrease in entropy for the Amazon reviews and Trivia-QA datasets, where for larger model sizes we don’t find improvements from RAIL approaches.
+In Table 4 we report RMSE on STSB dataset, MAE on Amazon reviews dataset, and $F_{1}$ metrics on Trivia-QA dataset from PaLM-2 models of varying size across multiple temperature values.
+We find improvements over baselines on STSB and Amazon reviews datasets for most temperatures.
+For Trivia-QA, we find improvements for XXS and S models for some temperatures, and for L, we don’t find a difference from our methods due to low entropy in the responses (see Table 7).
+In Table 10 we additionally report Pearson correlation metrics on STSB, confirming the results of RAIL improving over autoregressive inference.
+Lastly, in Table 9 we report cost weighted multi-class AUC-ROC with costs corresponding to the difference between the annotated labels: $|y_{1}-y_{2}|$.
+We find on both STSB and Amazon reviews datasets that the optimal decision rule (mean over the distribution) improves over the baselines.
 model STSB Amazon Trivia-QA
@@
 Table 8: RMSE as a function of the number of samples on STSB across PaLM-2 models of varying size. Results for temperature $T=0.25$.
-A.4 Related work
-LLMs for numerical tasks.
-Generative models have been successfuly applied to number prediction, where a number is generated token by token in an autoregressive manner.
-For example, Gruver et al.
-[2023] considered it in a zero-shot learning setup for time series prediction, and Liu and Low [2023], Yang et al.
-[2023] considered the autoregressive finetuning over numerical targets applied to arithmetic tasks.
-The importance of tokenizing the numerical targets into individual digits has been raised by previous works [Liu and Low, 2023, Yang et al., 2023].
-Finetuning approaches for target tasks alignment.
-Previous works considered approaches for aligning the models for target datasets.
-For example, soft prompts were finetuning on target datasets without loosing generalization to other tasks [Wang et al., 2023], and general finetuning was conducted on carefully tailored datasets for improved model robustness [Li et al., 2023].
-In our work, we focus on zero-shot setting where no fine-tuning is conducted.
-Finetuning approaches for numerical tasks.
-Autoregressive finetuning of LLMs on numerical tasks with CoT has been found effective [Liu and Low, 2023].
-One line of work for modeling predictive tasks with pre-trained Transformer based models is to add a regression head on top of the transformed/pooled encoded input tokens and finetune the resulting model on numerical targets using a regression loss.
-This is an approach which has been for encoder based models (e.g.
-Bert), and has also been applied to encoder-decoder (e.g.
-T5) models [Liu et al., 2022], and these approaches could be extended to decoder models too.
-In this work, we focus on the zero shot approaches, and so we leave training approaches for future work.
+model size greedy decode T=$0.25$ T=$0.5$ T=$1.0$
+argmax mean argmax mean argmax mean
+STSB XXS $0.797$ $0.755$ $0.882$ $0.714$ $0.890$ $0.632$ $0.889$
+XS $0.895$ $0.870$ $0.950$ $0.843$ $0.954$ $0.820$ $0.953$
+L $0.905$ $0.885$ $0.948$ $0.859$ $0.959$ $0.827$ $0.961$
+Amazon reviews XXS $0.87$ $0.894$ $0.925$ $0.866$ $0.94$ $0.788$ $0.942$
+XS $0.9$ $0.91$ $0.925$ $0.914$ $0.941$ $0.9$ $0.958$
+L $0.925$ $0.922$ $0.951$ $0.906$ $0.962$ $0.837$ $0.964$
+Table 9: Cost-weighted multi-partite AUC metrics on STSB and Amazon datasets (the higher the better).
+RAIL methods improve over the baselines.
+See §3.4 for the definition of AUC we use.
+We assume costs to correspond to the difference between the annotated labels: $|y_{1}-y_{2}|$.
+model greedy decode T=$0.25$ T=$0.5$ T=$1.0$
+argmax mean argmax mean argmax mean
+PaLM-2 XXS $0.767$ 0.738 $\mathbf{0.790}$ 0.670 $\mathbf{0.790}$ 0.544 $0.786$
+PaLM-2 XS $0.898$ 0.878 $\mathbf{0.915}$ 0.852 $0.913$ 0.821 $0.910$
+PaLM-2 L $0.909$ 0.893 $0.920$ 0.881 $0.922$ 0.860 $\mathbf{0.923}$
+Table 10: Pearson correlation metrics on STSB. RAIL methods improve over the baselines.
+In Table 8, we report the impact of the number of samples on the results.
+We note that there is an improvement in the results with the increase in the number of samples, however beyond 8 samples there is a diminishing improvement in practice.
+On STSB with temperature $\frac{1}{4}$, even with as few as two samples, our method starts to show improvements over greedy decoding.
+In Table 11, we report results for PaLM-2 models for RMSE on STSB when tuning the temperature parameter using a held-out set.
+Appendix E Why does sampling outperform enumeration?
+In this section we explicate why sampling can outperform enumeration. For easier reference, we first summarize what sampling and enumeration specifically mean:
+  • with the sampling strategy, we evaluate the average metric in (4) using $K$ samples drawn from the predictive distribution through temperature sampling.
+  • with the enumeration strategy, we score $K$ fixed targets from a uniform grid $G$, and replace the average metric in equation (4) with the estimate
+    $\sum_{g\in G}p(g)\cdot m(y,g)/\sum_{g\in G}p(g).$
+Now, a possible reason sampling performs better than enumeration can be that sampling is able to better explore the high density regions of the output probability space.
+For example, if the predictive distribution is concentrated in a tiny region of the output space, with the sampling strategy, most of the samples we use to estimate the optimal solution will be from this region.
+In contrast, with the enumeration strategy, most of the enumerated outputs will be from outside this region, and may prove not useful for estimating the optimal solution.
+For illustrative purposes, consider an extreme example for the STSB regression setup (where the output is a real number in $[0,5]$).
+Suppose the predictive distribution is a mixed probability distribution whose density is concentrated in a narrow region centered at $0.7$, and is near-uniform on all other targets.
+Since our enumeration strategy only considers the grid points $G=\{0.0,0.5,1.0,\ldots,5.0\}$, due to uniform probabilities over all these values it outputs:
+$\sum_{g\in G}p(g)\cdot g/\sum_{g\in G}p(g)=2.5.$
+With the sampling approach, all K samples will be drawn with high probability from the vicinity of $0.7$, and so, its output is:
+$\frac{1}{K}\sum_{k}\hat{y}_{k}\approx 0.7.$
+We would also like to note that both scoring and sampling improve over baselines, showing that both alternatives can make good use of the $\hat{p}(.|x)$.
+We also note that p may not be perfectly approximated by $\hat{p}$ due to various reasons, including the optimization, capacity, limited fine-tuning data and objectives used (e.g.
+label smoothing used in the pre-training objective).
+Appendix F Computational complexity of sampling-based RAIL
+Note that sampling can be done efficiently by caching the Transformer activations for the input prefix when generating different targets.
+In practice, when the prefix is long compared to the generated targets, a forward pass for the prefix tends to take most of the compute time.
+Note that is the case for scoring and regression tasks (the focus of our work), since the target score can be just a few tokens length, whereas the prefix can be long as it contains the input text.
+Moreover, we generate multiple samples simultaneously, and so, we do not incur a higher cost from generating multiple targets.
+model size greedy decode RAIL
+argmax mean
+XXS 1.047$\pm$0.004 1.447$\pm$0.007 0.967$\pm$0.004
+S 0.683$\pm$0.002 1.017$\pm$0.005 0.639$\pm$0.003
+L 0.628$\pm$0.003 0.988$\pm$0.004 0.578$\pm$0.002
+Table 11: Comparison of inference strategies on PaLM-2 models for RMSE on STSB when tuning the temperature on a held-out set.
+We draw $16$ samples.
+We use $\frac{1}{3}$ of the evaluation set for selecting the temperature from $\{0.25,0.5,0.75,1,2.5,5,7.5\}$, and use the remaining $\frac{2}{3}$ of the evaluation set for evaluation.
+We draw 10 random splits to obtain 95% confidence intervals.
+We confirm that the improvements that RAIL offers over baselines are indeed significant, and that when tuning the temperatures on the held-out set, the improvements hold.